Section 11

Mathematical Physics

In case of several authors, Invited Speakers are marked with a ^{∗}.

Eugene Bogomolny: Spectral Statistics . . . . . . . . . . . . . . . . . . . . . . . . . . . III 99

Detlev Buchholz: Scaling Algebras in Local Relativistic Quantum Physics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . III 109

J. T. Chayes: Finite-Size Scaling in Percolation . . . . . . . . . . . . . . . . . . . . . III 113
P. Collet: Extended Dynamical Systems . . . . . . . . . . . . . . . . . . . . . . . . . . . III 123
Robbert Dijkgraaf: The Mathematics of Fivebranes . . . . . . . . . . . . . . III 133

Antonio Giorgilli: On the Problem of Stability for Near to Integrable Hamiltonian Systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . III 143

Gian Michele Graf: Stability of Matter in Classical and Quantized Fields . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . III 153

Alexander Berkovich and Barry M. McCoy^{∗}:

Rogers-Ramanujan Identities: A Century of Progress from

Mathematics to Physics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . III 163

Roberto H. Schonmann: Metastability and the Ising Model . . . . . . . III 173

Feodor A. Smirnov: Space of Local Fields in Integrable Field Theory and Deformed Abelian Differentials . . . . . . . . . . . . . . . . . . . . . . . . III 183

Horng-Tzer Yau: Scaling Limit of Particle Systems, Incompressible Navier-Stokes Equation and Boltzmann Equation . . . . . . . . . . . . . . . . . . III 193

Documenta Mathematica · Extra Volume ICM 1998 · III ·

98

Documenta Mathematica · Extra Volume ICM 1998 · III ·